### v1

# Neutrino oscillations originate from virtual excitation of Z bosons

Abstract
To account for neutrino oscillations, it is postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation. However, I find that the energy conservation law requires these mass eigenstates, if they exist, to be entangled with distinct joint energy eigenstates of the other particles produced by the same weak interaction as the neutrino. This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which are responsible for flavor oscillations under the aforementioned postulation. I reveal that the neutrino oscillations actually originate from virtual excitation of the Z bosonic field diffusing over the space. During the propagation, the neutrino can continually excite and then immediately re-absorb a virtual Z boson. This virtual bosonic excitation produces a backaction on the neutrino, enabling it to oscillate among three flavors. When the neutrino propagates in matter, its behavior is determined by the competition between the coherent flavor transformation and decoherence effect resulting from scatterings.
In the electro-weak unification, the electromagnetic and weak interactions are described by a lepton model, where the Lagrangian has an exact symmetry, which is broken by the vacuum [1]. Within this framework, the intermediate bosons acquire their masses by coupling to a complex scalar field, which has a nonvanishing vacuum expectation value and pervades the space [2-4]. As a consequence, the weak interaction only occurs within a short range. All fermions except the neutrino pick up mass by the Yukawa interaction, which breaks the chiral symmetry. On the other hand, all neutrinos are left-handed, without right-handed counterparts, which makes the Yukawa interaction impossible.
The discovery of neutrino flavor transformations represents one of the most striking advancements of physical science achieved during the past few decades [5-29]. Such phenomena are commonly believed to show that neutrinos have nonvanishing mass [8-31], which is beyond the standard model. To account for these phenomena, it is assumed that there exist three distinct neutrino mass eigenstates. Each flavor eigenstate corresponds to a quantum-mechanical superposition of these mass eigenstates, and vice versa [32]. During the propagation of the neutrino, the probability amplitude associated with each mass eigenstate accumulates a phase that is approximately proportional to the square of the corresponding mass. The time-evolving phase differences among these probability amplitudes give arise to flavor oscillations, as a consequence of their quantum interference. Such an interpretation, however, overlooks the crucial question: Does nature allow neutrinos produced by weak interactions to be in such superposition states? As all neutrinos that were observed to exhibit flavor transformations had been generated through weak interactions, only when the answer of this question is affirmative can the aforementioned interpretation be reasonable. However, I find the answer is negative for the following reason. If the neutrino (or antineutrino) produced by a weak interaction has three distinct mass eigenstates ($\left|\nu_{i}\right\rangle$), these eigenstates should be correlated with different joint states ($\left|\psi_{i}\right\rangle$) of the other particles (accompanying particles) produced by the same interaction, e.g., the joint states for the proton and electron produced by a $\beta$ decay. The energy differences of these joint states need to compensate for those of the neutrino mass eigenstates in order to conserve the total energy. As detailed below, this correlation would destroy quantum interference effects among the neutrino mass eigenstates.
Suppose that the entire system just before participating in the weak interaction (e.g., the neutron just before making a $\beta$ decay) has a definite total energy $E_{T}$. Then the entire state of the neutrino and accompanying particles produced by such an interaction can be written as $\mathop{\displaystyle\sum}\limits_{i}C_{i}\left|\nu_{i}\right\rangle\left|\psi_{i}\right\rangle$. The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}-m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$. Only with such an arrangement can the final system has the same definite total energy, $E_{T}$, as the initial one, which ensures the process to comply with the energy conservation law. Under the assumption $m_{1}\neq m_{2}\neq m_{3}$, we have $E_{1}\neq E_{2}\neq E_{3}$, which implies that $\left|\psi_{1}\right\rangle$, $\left|\psi_{2}\right\rangle$, and $\left|\psi_{3}\right\rangle$ correspond to three energy eigenstates with different eigenvalues for the subsystem combined by the accompanying particles, thereby being orthogonal to each other. Due to this orthogonality, when the freedom degrees of the accompanying particles are traced out, the neutrino is left in a classically mixed state, described by the density operator, $\rho=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}\right|^{2}\left|\nu_{i}\right\rangle\left\langle\nu_{i}\right|$. Consequently, there does not exist any quantum coherence among the mass eigenstates, so that their mass differences cannot result in any flavor oscillation.
Moreover, if neutrino flavor eigenstates are defined as superpositions of mass eigenstates, it is impossible to creat a definite flavor by any weak interaction. This is due to the fact that, for the classical mixture $\rho$, the populations of the three flavor eigenstates ($\left|\nu_{\alpha}\right\rangle$) are all nonzero, given by $P_{\alpha}=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}G_{i\alpha}\right|^{2}$, where $\left|\nu_{i}\right\rangle=\mathop{\displaystyle\sum}\limits_{i}G_{i\alpha}\left|\nu_{\alpha}\right\rangle$. Therefore, the postulation that the flavor eigenstates are inconsistent with mass eigenstates not only fails to explain neutrino oscillations, but also leads to the wrong conclusion that the neutrino produced by any weak interaction has mixed flavors.
Here I reveal that neutrino oscillations actually originate from the neutrino’s interaction with the Z bosonic field, which diffuses over the whole space. As a consequence of vacuum fluctuation, the neutrino continually emits and re-absorbs virtual Z bosons, which produces a backaction on the neutrino. If the Z bosonic field can connect different neutrino flavors, such a backaction can transform these flavors into each other. When the neutrino propagates in matter, its behavior can be described by a master equation, featuring the competition between the coherent Hamiltonian dynamics and decoherence effect, caused by charged- and neutral-current scatterings. The observed flavor transformations for both the solar and atmospheric neutrinos can be well interpreted by such a master equation.
The physics underlying the neutrino flavor transformation can be described by the modified lepton model, where the Z bosonic field can connect different flavors of the neutrino field. The dynamics of the system combined by the neutrino field and Z bosonic field is governed by the Lagrangian [1,33]
${\cal L}_{n+Z}={\cal L}_{f}+{\cal L}_{m}+{\cal L}_{I}.$ (1)
${\cal L}_{f}$ is the free part without considering any coupling, which, to the second order, can be written as
${\cal L}_{f}=-\frac{1}{4}\left(\frac{\partial{\bf Z}_{\lambda}}{\partial x_{\nu}}-\frac{\partial{\bf Z}_{\nu}}{\partial x_{\lambda}}\right)^{2}-{\bf\phi}_{\alpha}^{\dagger}\gamma_{4}\gamma_{\lambda}\frac{\partial}{\partial x_{\lambda}}{\bf\phi}_{\alpha},$ (2)
where ${\bf\phi}_{\alpha}$ represents the field for the neutrino with flavor $\alpha$ ($\alpha=e,\mu,\tau$), and ${\bf Z}_{\lambda}$ denotes the Z bosonic field. ${\cal L}_{m}$ is the mass part, given by
${\cal L}_{m}=-\frac{1}{2}M_{Z}^{2}{\bf Z}_{\mu}^{2},$ (3)
where $M_{Z}$ is the mass of the $Z$ boson, gained by coupling to the Higgs field. The interaction between the neutrino field and the Z bosonic field is described by
${\cal L}_{I}=i\eta{\bf Z}_{\lambda}U_{\alpha,\beta}{\bf\phi}_{\alpha}^{\dagger}\gamma_{4}\gamma_{\lambda}(1+\gamma_{5}){\bf\phi}_{\beta},$ (4)
[…]
When $U_{\alpha,\beta}\neq 0$, the neutrino can transform between $\alpha$- and $\beta$-types of flavors by emitting and then immediately re-absorbing a virtual Z boson. To quantitatively describe thus-realized flavor transformation, we expand ${\bf Z}_{\lambda}$ in terms of the complete set of mode functions within a box of volume $V$ [33],
${\bf Z}_{T}$ $=$ $\mathop{\displaystyle\sum}\limits_{{\bf k}}\frac{1}{\sqrt{2\omega V}}{\bf a}_{T}{\bf(k)}e^{i{\bf k\cdot r}}+H.c.,$ (5)
${\bf Z}_{3}$ $=$ $\mathop{\displaystyle\sum}\limits_{{\bf k}}\frac{1}{\sqrt{2\omega V}}\frac{\omega}{M_{Z}}{\bf a}_{L}{\bf(k)}e^{i{\bf k\cdot r}}+H.c.,$
${\bf Z}_{4}$ $=$ $\mathop{\displaystyle\sum}\limits_{{\bf k}}\frac{i}{\sqrt{2\omega V}}\frac{k}{M_{Z}}{\bf a}_{L}{\bf(k)}e^{i{\bf k\cdot r}}+H.c.,$
where $\omega=\sqrt{M_{Z}^{2}+k^{2}}$, $T=1,2$ denotes the transverse polarization degrees of freedom, and ${\bf a}_{L}({\bf k})$ and ${\bf a}_{T}({\bf k})$ represent the annihilation operators for the longitudinal and transverse polarizations with the wavevector ${\bf k}$. The neutrino field can be expressed as
${\bf\phi}_{\alpha}=\sqrt{\frac{1}{V}}\mathop{\displaystyle\sum}\limits_{{\bf k}}{\bf b}_{\alpha}({\bf k})e^{i{\bf k\cdot r}},$ (6)
where ${\bf b}_{\alpha}$ denote the annihilation operator of the neutrino field with the flavor $\alpha$. As we here focus on the left-handed neutrino, the right-handed anti-neutrino part is not included in ${\bf\phi}_{\alpha}$. Then the Hamiltonian for neutrino and the Z bosonic field is given by
${\cal H}_{n+Z}={\cal H}_{0}+{\cal H}_{I},$ (7)
Here ${\cal H}_{0}$ corresponds to the ”free” part without considering the neutrino-boson interaction, given by [33]
${\cal H}_{0}=\mathop{\displaystyle\sum}\limits_{{\bf k}}k{\bf b}_{\alpha}^{\dagger}({\bf k}){\bf b}_{\alpha}({\bf k})+\mathop{\displaystyle\sum}\limits_{{\bf k}}\omega\left[{\bf a}_{L}^{\dagger}{\bf(k)a}_{L}{\bf(k)}+\mathop{\displaystyle\sum}\limits_{T=1,2}{\bf a}_{T}^{\dagger}{\bf(k)a}_{T}{\bf(k)}\right].$ (8)
The interaction part ${\cal H}_{I}$ is
[…]
$=$ $-\frac{i\eta}{\sqrt{2\omega V}}U_{\alpha,\beta}\mathop{\displaystyle\sum}\limits_{{\bf k,k}^{\prime}}{\bf X}_{\lambda}{\bf b}_{\alpha}^{\dagger}({\bf k})\gamma_{4}\gamma_{\lambda}(1+\gamma_{5}){\bf b}_{\beta}({\bf k}^{\prime}),$
where
${\bf X}_{T}$ $=$ ${\bf a}_{T}{\bf(k-k^{\prime})}+{\bf a}_{T}^{\dagger}{\bf(k}^{\prime}{\bf-k)},$ (10)
${\bf X}_{3}$ $=$ $\frac{\omega^{\prime}}{M_{Z}}\left[{\bf a}_{L}{\bf(k-k^{\prime})}+\omega^{\prime}{\bf a}_{T}^{\dagger}{\bf(k}^{\prime}{\bf-k)}\right],$
${\bf X}_{4}$ $=$ $\frac{i\left|{\bf k-k^{\prime}}\right|}{M_{Z}}\left[{\bf a}_{L}{\bf(k-k^{\prime})}-{\bf a}_{T}^{\dagger}{\bf(k}^{\prime}{\bf-k)}\right],$
with $\omega^{\prime}=\sqrt{M_{Z}^{2}+\left|{\bf k-k^{\prime}}\right|^{2}}$.
As the energy of the neutrino is much smaller than $M_{Z}$, the neutrino cannot emit any real Z gauge boson. However, the virtual excitation of the Z boson can induce effective couplings among different neutrino flavors. Suppose that the system is initially in the state $\left|\Psi_{i}\right\rangle=\left|\nu_{\alpha}({\bf p})\right\rangle\left|0\right\rangle$, where ${\bf p}$ denotes the momentum of the neutrino, and $\left|0\right\rangle$ is the vacuum state of the Z boson field. This initial state can be coupled to the final state $\left|\Psi_{f}\right\rangle=\left|\nu_{\beta}({\bf p}),0\right\rangle$ through the intermediate states $\left|\Psi_{m}^{j}({\bf k})\right\rangle={\bf a}_{j}^{\dagger}({\bf p}-{\bf k})\left|\nu_{\gamma}({\bf k}),0\right\rangle$, where $j=T,L$. Based on the second-order perturbation theory, ${\cal H}_{int}$ can be replaced by the effective Hamiltonian
${\cal H}_{eff}=\xi_{\alpha,\beta}\left|\nu_{\alpha}({\bf p}),0\right\rangle\left\langle\nu_{\beta}({\bf p}),0\right|,$ (11)
where
$\xi_{\alpha,\beta}$ $\simeq$ $-\mathop{\displaystyle\sum}\limits_{{\bf k}}\frac{\left\langle\Psi_{i}\right|{\cal H}_{int}\left|\Psi_{m}^{j}({\bf k})\right\rangle\left\langle\Psi_{m}^{j}({\bf k})\right|{\cal H}_{int}\left|\Psi_{f}\right\rangle}{M_{z}}$
$\simeq$ $-\mathop{\displaystyle\sum}\limits_{{\bf k}}\frac{6\eta^{2}U_{\alpha,\gamma}U_{\gamma,\beta}}{M_{z}^{2}V}$
$=$ $-\frac{3\eta^{2}}{4\pi^{3}M_{z}^{2}}\int d^{3}{\bf k}U_{\alpha,\gamma}U_{\gamma,\beta}.$
The integral divergence can be eliminated if $U_{\alpha,\gamma}$ contains a decaying factor, such as $e^{-\Gamma\left|{\bf k-p}\right|^{2}}$, which implies that the coupling between the initial and intermediate neutrino states exponentially decreases with the increasing momentum-phase distance. The result shows that the neutrino can change its flavor through virtual excitation of Z bosons, which can be understood as a backaction of the gauge field on the neutrino.
We now proceed to qualitatively interpret previous neutrino experiments based on this physical mechanism. Discarding the degrees of freedom of the Z bosonic field and performing the transformation $e^{i(p+\xi_{\tau,\tau})t}$, the neutrino dynamics during its propagation in the vacuum is governed by the Hamiltonian
${\cal H}$ $=$ $\Delta\left|\nu_{e}\right\rangle\left\langle\nu_{e}\right|+(\Delta+\delta)\left|\nu_{\mu}\right\rangle\left\langle\nu_{\mu}\right|$
$+(\xi_{e,\mu}\left|\nu_{e}\right\rangle\left\langle\nu_{\mu}\right|+\xi_{\tau,\mu}\left|\nu_{\tau}\right\rangle\left\langle\nu_{\mu}\right|$
$+\xi_{\tau,e}\left|\nu_{\tau}\right\rangle\left\langle\nu_{e}\right|+H.c.),$
where $\Delta=\xi_{e,e}-\xi_{\tau,\tau}$ and $\delta=\xi_{\mu,\mu}-\xi_{e,e}$. We here suppose all the Hamiltonian parameters are real. Under the condition $\delta,\xi_{e,\mu}\ll\xi_{\tau,\alpha}\ll\Delta$ with $\alpha=e,\mu$, the netrino evolution exhibits both fast and slow oscillations. For the short timescale $T\ll 1/\lambda$ with $\lambda=\xi_{e,\mu}+\xi_{\tau,e}\xi_{\tau,\mu}/\Delta$, the transition probability from the initial state $\left|\nu_{\alpha}\right\rangle$ ($\alpha=e,\mu$) to $\left|\nu_{\tau}\right\rangle$, caused by the fast oscillation, is approximately given by
$P_{f}(\left|\nu_{\alpha}\right\rangle\rightarrow\left|\nu_{\tau}\right\rangle)\simeq\frac{1}{2}A_{f}^{\alpha}[1-\cos(\Delta t)],$ (14)
where $A_{f}^{\alpha}=4\xi_{\tau,\alpha}^{2}/\Delta^{2}$. When $A_{s}^{\alpha}\ll 1$, such an oscillation has a small amplitude. The electron-antineutrino disappearance observed in the reactor neutrino experiments [19-21] can be described by this equation. The slow oscillation occur between $\left|\nu_{e}\right\rangle$ and $\left|\nu_{\mu}\right\rangle$. The resulting transition probability is well approximated by
$P_{s}(\left|\nu_{e}\right\rangle\longleftrightarrow\left|\nu_{\mu}\right\rangle)\simeq\frac{1}{2}A_{s}[1-\cos(\Omega_{l}t)].$ (15)
where $\Omega=\sqrt{4\lambda^{2}+\delta^{\prime 2}}$ and $A_{s}=\left|2\lambda/\Omega\right|^{2}$ with $\delta^{\prime}=\delta+(\xi_{\tau,\mu}^{2}-\xi_{\tau,e}^{2})/\Delta$. For $\lambda\gtrsim\left|\delta^{\prime}\right|$, we have $A_{s}\sim 1$, which qualitatively coincides with the KamLAND experimental result [22,23], although the quantitative dependences of the Hamiltonian parameters $\xi_{\alpha,\beta}$ on the neutrino energy are still unclear.
The problem of solar neutrino transformations is more complex. During the propagation of the neutrino from the core to surface of the sun, it may interact with other fermions through charged- or neutral-current scatterings [34-36], which would cause a decoherence effect, as interpreted below. Since the energy of the solar neutrino is much smaller than the masses of muon and tau, the reactions $\nu_{\mu/\tau}+e\rightarrow\mu/\tau+\nu_{e}$ cannot occur. Therefore, for the system combined by a neutrino and an electron the charged-current interaction can be described by the Lagrangian, $\frac{G}{\sqrt{2}}j_{\lambda}^{+}j_{\lambda}^{-}$, where $j_{\lambda}^{+}=(j_{\lambda}^{-})^{\dagger}=i{\bf\phi}_{e}^{+}\gamma_{4}\gamma_{\lambda}(1+\gamma_{5}){\bf\psi}_{e}$ with ${\bf\psi}_{e}$ denoting the electron field. As $j_{\lambda}^{+}$ ($j_{\lambda}^{-}$) transforms $e$ ($\nu_{e}$) into $\nu_{e}$ ($e$), this interaction swaps $e$ and $\nu_{e}$. For simplicity, we suppose that such a composite system is initially in the state $(u\left|\nu_{e}\right\rangle+v\left|\nu_{\mu}\right\rangle)\left|e\right\rangle$. Once the swapping event occurs, the system evolves into the entangled state $u\left|e\right\rangle\left|\nu_{e}\right\rangle+v\left|\nu_{\mu}\right\rangle\left|e\right\rangle$, where each particle is in a classical mixture of the neutrino- and electron-states when the other is traced out. In other words, after $\nu_{e}$-$e$ swapping the $\left|\nu_{e}\right\rangle$-state is transferred from the original neutrino to the original electron but the $\left|\nu_{\mu}\right\rangle$-state is not, so that they are carried by two different particles and thus cannot interfere with each other any more. Without quantum coherence, their combination corresponds to a classical mixture, described by the density operator $\left|u\right|^{2}\left|\nu_{e}\right\rangle\left\langle\nu_{e}\right|+\left|v\right|^{2}\left|\nu_{\mu}\right\rangle\left\langle\nu_{\mu}\right|$. The $\nu_{\alpha}$-$e$ elastic scattering process, resulting from the neutral-current interaction, also deteriorates the coherence since the $e$-type neutrino has a much higher probability to participate in such a process than the others [28]. When only the $e$-type neutrino elastically scatters an electron, $\left|\nu_{e}\right\rangle$ and $\left|\nu_{\mu}\right\rangle$ will be entangled with different momentum states, which would destroy the $\left|\nu_{e}\right\rangle$-$\left|\nu_{\mu}\right\rangle$ coherence. We note that such decoherence effects were not taken into account in previous investigations of neutrino propagation in matter [34-36]. Phenomenally, the competition between the coherent flavor coupling and incoherent scattering associated with $\left|\nu_{e}\right\rangle$ can be described by the master equation
$\frac{d\rho}{dt}=i\left[\rho,{\cal H}\right]+\frac{\gamma}{2}(2\sigma_{e}\rho\sigma_{e}-\sigma_{e}\rho-\rho\sigma_{e}),$ (16)
where $\sigma_{e}=\left|\nu_{e}\right\rangle\left\langle\nu_{e}\right|$. The decoherence rate $\gamma$ depends upon the neutrino energy and the properties of the matter. As the scattering cross-section scales with the neutrino energy, the high-energy neutrinos would have evolved into a steady state before reaching the solar surface, described by the density operator $\rho=\frac{1}{3}\mathop{\displaystyle\sum}\limits_{\alpha}\left|\nu_{\alpha}\right\rangle\left\langle\nu_{\alpha}\right|$, which remains invariant during the subsequent propagation. Consequently, the proportion of detected $\left|\nu_{e}\right\rangle\rightarrow\left|\nu_{e}\right\rangle$ events is $1/3$, as revealed in solar 8B neutrino experiments [7,12]. If the neutrino is free of such incoherent scatterings, the average fraction of the $\left|\nu_{e}\right\rangle$-outcomes after a very long time propagation is $1-\frac{1}{2}A_{s}$. For the intermediate case, the value is between these two extremes [37].
The atmospheric neutrinos originate from the pion decay and subsequent muon decay, which produce both the $\mu$- and $e$-type neutrinos with the numbers $N_{\mu}^{0}\simeq 2N_{e}^{0}$. According to the Super-Kamiokande experiments [17], when the Zenith angle ($\Theta$) ranges from $0^{\circ}$ to $80^{\circ}$, the disappearance probability of the $\mu$-type events for multi-GeV neutrinos exhibits slight oscillations, reflecting the comprehensive effect of the fast oscillation to $\left|\nu_{\tau}\right\rangle$ and the slow $\left|\nu_{e}\right\rangle\longleftrightarrow\left|\nu_{\mu}\right\rangle$ oscillation. Due to the imbalance between $N_{\mu}^{0}$ and $N_{e}^{0}$, for the number of $\mu$-type events ($N_{\mu}$) the gain from the $\left|\nu_{e}\right\rangle\rightarrow\left|\nu_{\mu}\right\rangle$ transition is less than the loss due to $\left|\nu_{\mu}\right\rangle\rightarrow\left|\nu_{e}\right\rangle$ transition. For $N_{e}$, the loss due to $\left|\nu_{e}\right\rangle\rightarrow\left|\nu_{\tau}\right\rangle$ transfer is partly compensated for by the net gain from the $\left|\nu_{e}\right\rangle\longleftrightarrow\left|\nu_{\mu}\right\rangle$ oscillation, so that its deficit is much smaller than that of $N_{\mu}$. With the extension of the propagation distance in the vacuum, the slow $\left|\nu_{e}\right\rangle\longleftrightarrow\left|\nu_{\mu}\right\rangle$ oscillation plays an increasingly important role. Within the regime $80^{\circ}<\Theta<90^{\circ}$, the number of net $\left|\nu_{\mu}\right\rangle\rightarrow\left|\nu_{e}\right\rangle$ transfer events surpasses that of $\left|\nu_{e}\right\rangle\rightarrow\left|\nu_{\tau}\right\rangle$ ones. When $\Theta>90^{\circ}$, the neutrino evolves under the competition between the coherent Hamiltonian dynamics and decoherence effect due to interactions with matter of the Earth. For multi-GeV neutrinos, the reactions $\nu_{\mu/\tau}+e\rightarrow\mu/\tau+\nu_{e}$ are allowed, but their occurrence probabilities would be much smaller than that for $\nu_{e}+e\rightarrow e+\nu_{e}$, since the transformations from $\nu_{\mu/\tau}$ into $\mu$ need to cost much more kinetic energy than the $\nu_{e}\rightarrow e$ transformation. As the sensitivities of the charged- and neutral-current reactions strongly depend upon the type of the neutrino, they will deteriorate the coherence between the neutrino flavors. For $90^{\circ}<\Theta<100^{\circ}$, the neutrino retains some quantum coherence after passing through the Earth, exhibiting slight oscillations. With the increase of $\Theta$, the slow $\left|\nu_{e}\right\rangle\longleftrightarrow\left|\nu_{\mu}\right\rangle$ oscillation has a higher probability of being interrupted by scattering events, which prevents monotonous increase of the net gain of $N_{e}$ from such an oscillation. When $100^{\circ}<\Theta<180^{\circ}$, the neutrino has approximately evolved to the maximally mixed state $\rho$ after crossing the Earth, so that the three flavors roughly have the same population, independent of the initial state.
As our theoretical model allows occurrence of oscillations even if the neutrino is massless, one may argue that the state of a massless particle cannot be changed during its propagation in the vacuum. This is the case if the particle remains unperturbed, as exemplified by the free propagation of a photon. However, the photonic state can be changed by some scattering process. For example, a photon can transform between modes $a$ and $b$ with the assistance of $\Lambda$-type three-level atoms, which have an excited state $\left|e\right\rangle$ and two ground states $\left|g\right\rangle$ and $\left|f\right\rangle$ [38]. The transitions $\left|e\right\rangle\rightarrow\left|g\right\rangle$ and $\left|e\right\rangle\rightarrow\left|f\right\rangle$ are respectively coupled to these two modes with the same detuning. When the detuning is much larger than the coupling strengths, the photon would oscillate between these two modes through a Raman scattering process, mediated by virtual excitation of the atoms. The flavor oscillations of the neutrino bear some similarity to such a Raman process. As the neutrino is inevitably disturbed by the vacuum fluctuations of the gauge fields pervading the space, its propagation cannot be really free even in the vacuum. The backaction of the virtually excited Z gauge field on the neutrino is analogous to that of the virtually excited atoms on the photon in the Raman process. One may further ask: Does there exist any other mechanism that can mediate the neutrino flavor transformation? The answer is presumably negative for the following reason. The W and Z bosonic fields are the only elements that can interact with the neutrino during its propagation in the vacuum. As the neutrino flavor is defined by the charged current that is mediated by the W bosonic field, the vacuum fluctuation of the Z bosonic field seems to be the only mechanism that can change the neutrino flavor.
This work was supported by the National Natural Science Foundation of China under Grant No. 12274080.

### v2

@@
 However, I find that the energy conservation law requires these mass eigenstates, if they exist, to be entangled with distinct joint energy eigenstates of the other particles produced by the same weak interaction as the neutrino.
-This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which are responsible for flavor oscillations under the aforementioned postulation.
+This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which is responsible for flavor oscillations under the aforementioned postulation.
 I reveal that the neutrino oscillations actually originate from virtual excitation of the Z bosonic field diffusing over the space.
@@
 Then the entire state of the neutrino and accompanying particles produced by such an interaction can be written as $\mathop{\displaystyle\sum}\limits_{i}C_{i}\left|\nu_{i}\right\rangle\left|\psi_{i}\right\rangle$.
-The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}-m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$.
+The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}+m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$.
 Only with such an arrangement can the final system has the same definite total energy, $E_{T}$, as the initial one, which ensures the process to comply with the energy conservation law.
@@
 Consequently, there does not exist any quantum coherence among the mass eigenstates, so that their mass differences cannot result in any flavor oscillation.
-Moreover, if neutrino flavor eigenstates are defined as superpositions of mass eigenstates, it is impossible to creat a definite flavor by any weak interaction.
-This is due to the fact that, for the classical mixture $\rho$, the populations of the three flavor eigenstates ($\left|\nu_{\alpha}\right\rangle$) are all nonzero, given by $P_{\alpha}=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}G_{i\alpha}\right|^{2}$, where $\left|\nu_{i}\right\rangle=\mathop{\displaystyle\sum}\limits_{i}G_{i\alpha}\left|\nu_{\alpha}\right\rangle$.
-Therefore, the postulation that the flavor eigenstates are inconsistent with mass eigenstates not only fails to explain neutrino oscillations, but also leads to the wrong conclusion that the neutrino produced by any weak interaction has mixed flavors.
 Here I reveal that neutrino oscillations actually originate from the neutrino’s interaction with the Z bosonic field, which diffuses over the whole space.

### v3

@@
 Abstract
-To account for neutrino oscillations, it is postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation.
-However, I find that the energy conservation law requires these mass eigenstates, if they exist, to be entangled with distinct joint energy eigenstates of the other particles produced by the same weak interaction as the neutrino.
+To account for neutrino oscillations, it was postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation.
+However, for a neutrino produced by the decay of an unstable particle, these mass eigenstates, if they exist, would be entangled with distinct joint momentum eigenstates of the entire system, composed of all the particles produced by the decay.
 This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which is responsible for flavor oscillations under the aforementioned postulation.
@@
 The time-evolving phase differences among these probability amplitudes give arise to flavor oscillations, as a consequence of their quantum interference.
-Such an interpretation, however, overlooks the crucial question: Does nature allow neutrinos produced by weak interactions to be in such superposition states?
-As all neutrinos that were observed to exhibit flavor transformations had been generated through weak interactions, only when the answer of this question is affirmative can the aforementioned interpretation be reasonable.
-However, I find the answer is negative for the following reason.
-If the neutrino (or antineutrino) produced by a weak interaction has three distinct mass eigenstates ($\left|\nu_{i}\right\rangle$), these eigenstates should be correlated with different joint states ($\left|\psi_{i}\right\rangle$) of the other particles (accompanying particles) produced by the same interaction, e.g., the joint states for the proton and electron produced by a $\beta$ decay.
-The energy differences of these joint states need to compensate for those of the neutrino mass eigenstates in order to conserve the total energy.
-As detailed below, this correlation would destroy quantum interference effects among the neutrino mass eigenstates.
-Suppose that the entire system just before participating in the weak interaction (e.g., the neutron just before making a $\beta$ decay) has a definite total energy $E_{T}$.
-Then the entire state of the neutrino and accompanying particles produced by such an interaction can be written as $\mathop{\displaystyle\sum}\limits_{i}C_{i}\left|\nu_{i}\right\rangle\left|\psi_{i}\right\rangle$.
-The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}+m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$.
-Only with such an arrangement can the final system has the same definite total energy, $E_{T}$, as the initial one, which ensures the process to comply with the energy conservation law.
-Under the assumption $m_{1}\neq m_{2}\neq m_{3}$, we have $E_{1}\neq E_{2}\neq E_{3}$, which implies that $\left|\psi_{1}\right\rangle$, $\left|\psi_{2}\right\rangle$, and $\left|\psi_{3}\right\rangle$ correspond to three energy eigenstates with different eigenvalues for the subsystem combined by the accompanying particles, thereby being orthogonal to each other.
-Due to this orthogonality, when the freedom degrees of the accompanying particles are traced out, the neutrino is left in a classically mixed state, described by the density operator, $\rho=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}\right|^{2}\left|\nu_{i}\right\rangle\left\langle\nu_{i}\right|$.
-Consequently, there does not exist any quantum coherence among the mass eigenstates, so that their mass differences cannot result in any flavor oscillation.
+However, as has been proved [33,34], if the state of a neutrino produced from the decay of an unstable particle consists of different mass eigenstates, the mass degree of freedom of the neutrino is necessarily entangled with the momentum freedom degrees of the accompanying particles or/and that of the neutrino itself, irrespective of the momentum distribution of the unstable particle.
+Due to this entanglement, the neutrino is essentially in a classical mixture of these mass eigenstates when all the momentum freedom degrees are traced out.
+Consequently, the neutrino cannot has a definite flavor at the production, and the population of each flavor eigenstate cannot be changed during the propagation.
 Here I reveal that neutrino oscillations actually originate from the neutrino’s interaction with the Z bosonic field, which diffuses over the whole space.
@@
 As the neutrino flavor is defined by the charged current that is mediated by the W bosonic field, the vacuum fluctuation of the Z bosonic field seems to be the only mechanism that can change the neutrino flavor.
-This work was supported by the National Natural Science Foundation of China under Grant No. 12274080.
+Finally, it should be noted that the virtual excitation of Z bosons cannot lead to flavor transformations between different flavors of charged leptons owing to their large mass differences.